OpenStudy (anonymous):

Find three consecutive numbers such that the sum of one-fourth the first and one-fifth the second is five less than one-seventh the third. -10, -11 and -12 -12, -13 and -14 -14, -15 and -16

OpenStudy (anonymous):

\[1/4a+1/5b=1/3c-5\] a=a b=a+1 c=a+1+1 If i read right

OpenStudy (anonymous):

so would it be the 2nd one?

OpenStudy (zepp):

Okay, let the first term be a, second term be a+1, and third term be a+2, since they are consecutive Then, write our equation. \(\large \frac{1}{4}a+\frac{1}{5}(a+1)=\frac{1}{7}(a+2)\)

OpenStudy (anonymous):

What zepp said. my 1/3 should be 1/7

OpenStudy (zepp):

Oh sorry, \(\large \frac{1}{4}a+\frac{1}{5}(a+1)=\frac{1}{7}(a+2)-5\) Should be like that, I forgot the 5 less.

OpenStudy (zepp):

And solve for a.